40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 6 00:40:29 2009, Alberto, Update, LSC, Everything put back as it was 
    Reply  Mon Nov 9 15:11:02 2009, Alberto, Update, LSC, Everything put back as it was 
Message ID: 2217     Entry time: Mon Nov 9 15:11:02 2009     In reply to: 2189
Author: Alberto 
Type: Update 
Category: LSC 
Subject: Everything put back as it was 

Quote:

I disconnected the setup for the arm cavity TF measurement. I opened the scitation switch on the ISS medm screen. I reconnected the OMC ISS EXC cable to the breakout box on the floor.

The photodiode on the Y end is stilll connected.

Also the RFAm (whish is not disbaled anymore) still has a 50% beam splitter before it.

I'm also running the Align Full IFO script.

 

I removed the beam splitter and the PDA 255.

the beam path to the RFAM photodiode is clear again.

ELOG V3.1.3-